A mineshaft has an ore elevator hung from a single braided cable of diameter 2.5 cm. Young's modulus of the cable is 10×1010N/m2.When the cable is fully extended, the end of the cable is 400 m below the support.

How much does the fully extended cable stretch when 1500 kg of ore is loaded?

Express your answer with the appropriate units.
Answer #1

stress = F/A

= m*g/(pi*d^2/4)

= 1500*9.8/(pi*0.025^2/4)

= 2.994*10^7 N/m^2

we know, young's modulus, Y = stress/strain

strain = streess/Y

delta_L/L = stress/Y

delta_L = L*stress/Y

= 400*2.994*10^7/(10*10^10)

= 0.119 m or 11.9 cm <<<<<<--------Answer